# Multifrequency Radio Observations of the Magnetar Swift J1818.0–1607

### v1

@@ 5 Discussion and Conclusions
Swift J1818.0–1607 possessed a steep, negative radio spectral index ($\alpha<-1.8$) for the first few months post-outburst (Maan & van Leeuwen 2020; Lower et al. 2020), and was not detectable at or above 8 GHz (Gajjar et al. 2020; Lower & Shannon 2020). Four months post-outburst, a significant spectral flattening was detected, with a spectral index below 9 GHz of $\alpha=+0.3(2)$ (Majid et al., 2020). Follow-up observations yielded high-frequency detections of J1818 up to 151 GHz, implying an overall flattening of the spectral index with time (Torne et al., 2020). Longer-term followup found that in the months preceding our observations, the spectral index of J1818 between 2 and 8 GHz flattened over time (Huang et al. 2021; Bansal et al. 2023).
@@
Most radio magnetars have, at least temporarily, shown evidence of radio spectra which peak at gigahertz frequencies (Gigahertz Peaked Spectrum; GPS). These include XTE J1810–197 (Eie et al., 2021), SGR 1745–2900 (Pennucci et al., 2015), 1E 1547.1–5408, and PSR J1622–4950 (Kijak et al., 2017). GPS are also observed in a very small subset of rotation-powered radio pulsars, albeit with lower peak frequencies (0.6 to 2 GHz) than the bulk of the radio magnetar population (Jankowski et al., 2018). The prevailing explanation for these spectral peaks is an interaction between the radio emission and the ionized environments around the neutron star; it is also possible that the younger magnetars are situated in denser or hotter environments, leading to higher peak frequencies (Kijak et al. 2013, 2017).
@@
The mean flux at 35 GHz is higher than the 22-GHz flux (see Table 3), which appears counterintuitive given the relative S/N ratios of each integrated profile. At such high frequencies, weather and atmospheric effects could have enhanced the system noise at 35 GHz. Since the 35-GHz observation took place 54 days before the 22-GHz observation, it is possible that the overall flux of J1818 simply decayed in that time, or that the flux in each band varies on shorter timescales. The total spectrum shown in Figure 4 displays better continuity between the fluxes at 22 and 35 GHz, especially given the time span between observations.
Simultaneous, multi-frequency observations are necessary to determine whether spectral turnovers are present in this and other radio magnetars. If the 35-GHz flux were consistently higher than the 22-GHz flux across many observational epochs, it is possible that there is another spectral turnover and that J1818 once again exhibits a positive spectral index above 30 GHz. XTE J1810–197 exhibited similar spectral features, with a large increase in the radio flux above 150 GHz (Eie et al., 2021). The magnetars SGR 1745-1900 (Torne et al., 2017) and 1E 1547.0–5408 (Chu et al., 2021) have shown evidence of spectral turn-ups in the range of 20–80 GHz.
@@ 5.2 Pulse Profile Morphology
Across the frequency range of our observations, the integrated profiles consisted of a narrow, consistent central component (II), flanked by more sporadic components (I and III) comprised of narrow sub-pulses. The outer components tended to have flatter spectra and higher levels of modulation in time, similar to the emission properties observed in radio magnetars. Furthermore, the majority of pulses at 9 GHz with the sign of the circular polarization reversed originated from the phase range of component III only.
Early observations of J1818 found that its integrated profile typically consisted of a single peak with a high degree of linear polarization and a steep spectrum (Lower et al. 2020; Champion et al. 2020). This profile component was consistently detected across multiple epochs and frequencies, and its properties are similar to those we observe in Component II. Brighter, more sporadic pulses were occasionally detected from a postcursor component as early as five days post-outburst (Maan & van Leeuwen 2020; Champion et al. 2020), suggesting it is the same sporadic postcursor component that we observe as Component III.
We only detect the precursor component (I) with high significance at 13 and 22 GHz, the latest two observations in our dataset. 12 pulses with a significance level $\geq 4$ (as defined in Section 4.1) were detected from the phase range of Component I in the 9-GHz dataset. A low S/N precursor component is visible by eye in the 35 GHz profile, though it is offset in longitude from the position of the precursor component at other frequencies.
Emission from a precursor component was also detected by Lower et al. (2021) and Huang et al. (2021). The precursor component detected at 2.5 GHz in Lower et al. (2021) seemed to appear sometime between MJDs 59009 and 59047, and over the next two months shifted in pulse longitude until it nearly overlapped the main pulse component at the end of their observations on MJD 59128. Similar drifts in the pulse longitude of individual pulse components were also observed in the radio magnetar XTE J1810–197 (Levin et al., 2019). This component may have a sufficiently steep spectrum to not be detectable at lower frequencies, or the longitudinal drift may have reversed direction between observations.
@@ 5.2.2 Mode Switching
Mode switching is often explained by changes in the large-scale configuration of the pulsar magnetosphere (Timokhin, 2010). Given our lack of detection of any kind of mode switching in our observations, it is possible that J1818 settled into a more stable emission state roughly five months post-outburst.
@@ 6 Conclusions
The integrated pulse profile of J1818 was also remarkably stable across frequencies, consisting of up to three pulse components. The central component (Component II) tends to be narrower and displays less flux variability over time when compared to the other two components. The postcursor (Component III) in particular is composed largely of narrow, bright subpulses which fall within the larger pulse envelope, reminiscient of the highly variable radio bursts from magnetars. At most frequencies, the outer profile components have flatter spectral indices and higher flux modulation indices than the central component.
@@
E.F.L. and M.A.M. are supported by NSF award AST-2009425. M.A.M. is also supported by NSF Physics Frontiers Center award PHYS-2020265. The Green Bank Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc.

### v2

@@ 5 Discussion and Conclusions
Swift J1818.0–1607 possessed a steep, negative radio spectral index ($\alpha<-1.8$) for the first few months post-outburst (Maan & van Leeuwen 2020; Lower et al. 2020; Champion et al. 2020), and was not detectable at or above 8 GHz (Gajjar et al. 2020; Lower & Shannon 2020). Four months post-outburst, a significant spectral flattening was detected, with a spectral index below 9 GHz of $\alpha=+0.3(2)$ (Majid et al., 2020). Follow-up observations yielded high-frequency detections of J1818 up to 154 GHz, implying an overall flattening of the spectral index with time (Torne et al., 2020). Longer-term followup found that in the months preceding our observations, the spectral index of J1818 between 2 and 8 GHz flattened over time (Huang et al. 2021; Bansal et al. 2023).
@@
Most radio magnetars have, at least temporarily, shown evidence of radio spectra which peak at gigahertz frequencies (Gigahertz Peaked Spectrum; GPS). These include XTE J1810–197 (Eie et al., 2021), SGR 1745–2900 (Pennucci et al., 2015), 1E 1547.1–5408, and PSR J1622–4950 (Kijak et al., 2013). GPS are also observed in a very small subset of rotation-powered radio pulsars, albeit with lower peak frequencies (0.6 to 2 GHz) than the bulk of the radio magnetar population (Jankowski et al., 2018). The prevailing explanation for these spectral peaks is an interaction between the radio emission and the ionized environments around the neutron star; it is also possible that the younger magnetars are situated in denser or hotter environments, leading to higher peak frequencies (Kijak et al. 2013, 2017).
@@ 5.2 Pulse Profile Morphology
Across the frequency range of our observations, the integrated profiles consisted of a narrow, consistent central component (II), flanked by more sporadic components (I and III) comprised of narrow sub-pulses. The outer components tended to be comprised of bright, narrow subpulses and had higher levels of flux modulation in time, similar to the emission properties observed in radio magnetars. Furthermore, the majority of pulses at 9 GHz with the sign of the circular polarization reversed originated from the phase range of component III only.
Early observations of J1818 found that its integrated profile typically consisted of a single peak with a high degree of linear polarization and a steep spectrum (Lower et al. 2020; Champion et al. 2020). This profile component was consistently detected across multiple epochs and frequencies, and its properties are similar to those we observe in Component II. Brighter, more sporadic pulses were occasionally detected from a trailing component as early as five days post-outburst (Maan & van Leeuwen 2020; Champion et al. 2020), suggesting it is the same sporadic trailing component that we observe as Component III.
We only detect the leading component (I) with high significance at 13 and 22 GHz, the latest two observations in our dataset. 12 pulses with a significance level $\geq 4$ (as defined in Section 4.1) were detected from the phase range of Component I in the 9-GHz dataset. A low S/N leading component is visible by eye in the 35 GHz profile, though it is offset in longitude from the position of the leading component at other frequencies.
Emission from a leading component was also detected by Lower et al. (2021) and Huang et al. (2021). The leading component detected at 2.5 GHz in Lower et al. (2021) seemed to appear sometime between MJDs 59009 and 59047, and over the next two months shifted in pulse longitude until it nearly overlapped the main pulse component at the end of their observations on MJD 59128. Similar drifts in the pulse longitude of individual pulse components were also observed in the radio magnetar XTE J1810–197 (Levin et al., 2019). This component may have a sufficiently steep spectrum to not be detectable at lower frequencies, or the longitudinal drift may have reversed direction between observations.
@@ 5.2.2 Mode Switching
High-cadence monitoring of J1818 at 1.4 and 6 GHz revealed that J1818 displayed three distinct emission modes during the course of our observing campaign, which provide important context for these results (Rajwade et al. 2022; Fisher et al. 2024). Our 6-GHz observation on MJD 59092 occurred during Mode I (as labeled by Rajwade et al. (2022)), when the integrated pulse profile at 1.4 and 6 GHz was single-peaked. We observe similar behavior in our 6 GHz profile, although we observe a low-intensity trailing component.
The first mode change occurred at around MJD 59100, after which the trailing component appeared, and the flux of the main component became more variable. The rest of our observations took place during Mode II, except for our observation at 13 GHz which took roughly 15 days after the switch to Mode III.
The Mode II observed by Rajwade et al. (2022) is marked by the narrow main component followed by a broader trailing component; the leading component is barely visible in their 6 GHz profile and not at all visible in their 1.4 GHz profile. In our observations taken during this mode, the leading component is barely visible at 9 GHz, but more prominent at 22 GHz. The mode change to Mode III took place eleven days after our 22-GHz observation according to Fisher et al. (2024), after which point the leading component became much more prominent. Mode changing is usually quasi-instantaneous, and the daily observing cadence of Rajwade et al. (2022) has localized the epoch of the mode change to within one day, so we cannot relate the appearance of this component with the reported mode change. The profile variability maps shown in Fisher et al. (2024) reveal that, the emission from the precursor component was weaker than average before MJD $\sim$59206, and it is not at all visible at 1.4 GHz in the integrated profile taken just a day before our 22 GHz observation. The appearance of the leading component in our integrated profiles before the reported mode change indicates the frequency- dependent evolution of this component, independent of the mode changing phenomenon.
@@ 6 Conclusions
The integrated pulse profile of J1818 was also remarkably stable across frequencies, consisting of up to three pulse components. The central component (Component II) tends to be narrower and displays less flux variability over time when compared to the other two components. The trailing component (Component III) in particular is composed largely of narrow, bright subpulses which fall within the larger pulse envelope, reminiscient of the highly variable radio bursts from magnetars. At most frequencies, the outer profile components have flatter spectral indices and higher flux modulation indices than the central component.
@@
We thank the anonymous reviewer for their comments which greatly improved the manuscript. E.F.L. and M.A.M. are supported by NSF award AST-2009425. M.A.M. is also supported by NSF Physics Frontiers Center award PHYS-2020265. The Green Bank Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc.